# BayesianFitForecast: A User-Friendly R Toolbox for Parameter Estimation and Forecasting with Ordinary Differential Equations

### v2

@@ 3.8.2 Run_analyzeResults
By leveraging the BayesianFitForecast library through CRAN, users can easily integrate Bayesian fitting and forecasting workflows into their R environment. Additionally, we provide an R Shiny App interface to facilitate use of the package in a graphical interface.
@@ 3.9.1 Installing and loading the BayesianFitForecast App
Prior to running the BayesianFitForecast R-Shiny application, the most recent version of R and RStudio must be downloaded. A system-specific package compiler must also be loaded (i.e., Rtools, XCode). Additional details can be found in the ”package_builder_instructions” document available within the Shiny App source folder on the GitHub page (https://github.com/gchowell/BayesianFitForecast/tree/main/Shiny%20app%20(Beta%20version)). Once the necessary software is installed, users can then:
@@
Required Software R (¿= 4.3), R-Studio (2024.09.0 Build 375), System-Specific Package Compiler
@@ 3.9.2 Dashboard Structure
The dashboard utilizes the same input data structure and naming scheme described in Section 2.4.14. To load the data into the dashboard, users first must navigate to the ”Data Options” menu on the sidebar, select the Browse button, and identify the file location within their personal computer. Once loaded, additional data specification parameters related to the process of interest included in the data (ex., ”Influenza”), its temporal resolution (ex., ”Days”), and the type of data included (ex., ”Cases”), will become available.
@@ A.2.1 Data Options
Figure 14: The ”Data Options” available within the sidebar menu of the BayesianFitForecast dashboard. The first step in utilizing the dashboard includes loading the input data (ex. SanFrancisco.xlsx) and specifying the process of interest (ex. <sanfranscio>), temporal resolution of the data (ex. <Days>), and label indicating the type of included time series (ex. <Cases>). The Process of Interest, Temporal Resolution, and Time Series Type # parameters are free-text inputs that are not required to follow any specific format.
@@ A.2.2 Forecasting Options
Figure 15: The ”Forecasting Options” available within the sidebar menu of the BayesianFitForecast dashboard. The second step in utilizing the dashboard includes specifying the calibration period length and forecasting horizon. The Calibration Period Length parameter is an auto-populated drop-down menu, and the Forecasting Horizon parameter is a numeric input. For this tutorial, we set the calibration period length to 17 and the forecasting horizon to 10.
@@ A.2.3 Base Model Options
The ”Base Model Options” menu located within the sidebar contains the base controls for the state variable, parameter, and composite equation specifications (Fig. 16). The Error Structure option is a drop-down menu containing the three available error structures: (1) Negative Binomial, (2) Normal, and (3) Poisson. Only one error structure can be selected at a time. As discussed in the main text, additional parameters are required to be estimated when the Negative Binomial or Normal error structures are selected. To accommodate this requirement, when either error structure is selected, the corresponding symbol for the required parameters will appear under the “Required Parameters” header on the “Parameter Specifications” tab. Additional details regarding prior selection for the required parameters can be found below. For our example, we selected the <Poisson> error structure option.
Figure 16: The ”Base Model” options available within the last menu of the sidebar. The figure illustrates the base model options available within the dashboard’s sidebar. Prior to specifying the state variable, parameters, and system of ordinary differential equations (ODE), users must select the error structure of interest (Error Structure) via the drop-down menu, and specify the number of desired state variables (# State Variables), parameters (# Parameters), composite variables (# Composite Variables), number of MCMC steps (Number of MCMC Steps), and the number of chains to include (Number of MCMC Chains) using the corresponding numeric input. For this tutorial, we indicated we wanted five state variables, four unique parameters, and two composite variables. We also employed 1000 MCMC steps and two MCMC chains.

### v3

@@ 3.9.2 Run_analyzeResults
By leveraging the BayesianFitForecast library through CRAN, users can easily integrate Bayesian fitting and forecasting workflows into their R environment. Additionally, we provide an R Shiny App interface that offers a graphical dashboard for the BayesianFitForecast workflow. This application enables users to upload data, configure model parameters, specify priors and ODE structures, and view diagnostic and forecasting outputs without writing code. The Shiny app mirrors the core functionality of the R scripts, allowing consistent and reproducible results through a more accessible interface.
@@ 3.10.1 Installing and loading the BayesianFitForecast App
Prior to running the BayesianFitForecast R-Shiny application, the most recent version of R and RStudio must be downloaded. A system-specific package compiler must also be loaded (i.e., Rtools, XCode). Additional details can be found in the "package_builder_instructions" document available within the Shiny App source folder on the GitHub page (https://github.com/gchowell/BayesianFitForecast/tree/main/Shiny%20app%20(Beta%20version)). Once the necessary software is installed, users can then:
@@
Required Software R (>= 4.3), R-Studio (2024.09.0 Build 375), System-Specific Package Compiler
@@ 3.10.2 Dashboard Structure
The dashboard utilizes the same input data structure and naming scheme described in Section 2.4.15. To load the data into the dashboard, users first must navigate to the "Data Options" menu on the sidebar, select the Browse button, and identify the file location within their personal computer. Once loaded, additional data specification parameters related to the process of interest included in the data (ex., "Influenza"), its temporal resolution (ex., "Days"), and the type of data included (ex., "Cases"), will become available.
@@ 4 Conclusions
Availability and requirements
{unenumerate}
Project name: BayesianFitForecast
Project home page: https://github.com/gchowell/BayesianFitForecast
Operating system(s): Platform independent
Programming language: R
Other requirements: R, RStudio, Java, and Rtools (or equivalent package builder)
List of abbreviations
{unenumerate}
ODEs: Ordinary Differential Equations
MCMC: Markov chain Monte Carlo
HMC: Hamiltonian Monte Carlo
NUTS: No U-Turn Sampler
DIC: Deviance Information Criterion
AIC: Akaike Information Criterion
PI: Prediction Interval
CI: Confidence Interval
MAE: Mean Absolute Error
MSE: Mean Squared Error
WIS: Weighted Interval Scores
IS: Interval Scores
SEIR: Susceptible-Exposed-Infected-Recovered model
SEIUR: Susceptible-Exposed-Infected-Under-reported model
SEIRD: Susceptible-Exposed-Infected-Recovered-Death model
Declarations
{unenumerate}
Ethics approval and consent to participate: Not applicable.
Consent for publication: Not applicable.
Availability of data and materials: This program and all associated files and data mentioned within the manuscript are publicly available at https://github.com/gchowell/BayesianFitForecast.
Competing interests: Not applicable
Funding: HK and AB are supported by a 2CI Fellowship from Georgia State University. G.C. is partially supported by NSF grants 2125246 and 2026797. The funding bodies played no role in the design of the study and collection, analysis, and interpretation of data and in writing the manuscript.
Authors’ contributions: H.K., R.L., and G.C. conceived and developed the first version of the toolbox, with H.K. responsible for coding. H.K., R.L., and G.C. wrote the first draft of the tutorial. H.K., A.B., R.L., and G.C. contributed to writing and reviewing subsequent drafts of the tutorial. A.B. produced the tutorial video. A.B., H.K., and G.C. generated the Shiny App, with A.B. responsible for the UI and related coding. A.B. wrote the first draft related to the Shiny App and Appendix, and H.K. and G.C. contributed to subsequent drafts.
Acknowledgements: Not applicable
@@ A.2.1 Data Options
Figure 16: The "Data Options" available within the sidebar menu of the BayesianFitForecast dashboard. The first step in utilizing the dashboard includes loading the input data (ex. SanFrancisco.xlsx) and specifying the process of interest (ex. <sanfranscio>), temporal resolution of the data (ex. <Days>), and label indicating the type of included time series (ex. <Cases>). The Process of Interest, Temporal Resolution, and Time Series Type # parameters are free-text inputs that are not required to follow any specific format.
@@ A.2.2 Forecasting Options
Figure 17: The "Forecasting Options" available within the sidebar menu of the BayesianFitForecast dashboard. The second step in utilizing the dashboard includes specifying the calibration period length and forecasting horizon. The Calibration Period Length parameter is an auto-populated drop-down menu, and the Forecasting Horizon parameter is a numeric input. For this tutorial, we set the calibration period length to 17 and the forecasting horizon to 10.
@@ A.2.3 Base Model Options
The "Base Model Options" menu located within the sidebar contains the base controls for the state variable, parameter, and composite equation specifications (Fig. 18). The Error Structure option is a drop-down menu containing the three available error structures: (1) Negative Binomial, (2) Normal, and (3) Poisson. Only one error structure can be selected at a time. As discussed in the main text, additional parameters are required to be estimated when the Negative Binomial or Normal error structures are selected. To accommodate this requirement, when either error structure is selected, the corresponding symbol for the required parameters will appear under the “Required Parameters” header on the “Parameter Specifications” tab. Additional details regarding prior selection for the required parameters can be found below. For our example, we selected the <Poisson> error structure option.
Figure 18: The "Base Model" options available within the last menu of the sidebar. The figure illustrates the base model options available within the dashboard’s sidebar. Prior to specifying the state variable, parameters, and system of ordinary differential equations (ODE), users must select the error structure of interest (Error Structure) via the drop-down menu, and specify the number of desired state variables (# State Variables), parameters (# Parameters), composite variables (# Composite Variables), number of MCMC steps (Number of MCMC Steps), and the number of chains to include (Number of MCMC Chains) using the corresponding numeric input. For this tutorial, we indicated we wanted five state variables, four unique parameters, and two composite variables. We also employed 1000 MCMC steps and two MCMC chains.